A charging RC circuit controls the intermittent windshield wipers in a car. The emf is 12.0 V. The wipers are triggered when the voltage across the 110 µF capacitor reaches 10.0 V; then the capacitor is quickly discharged (through a much smaller resistor) and the cycle repeats. What resistance should be used in the charging circuit if the wipers are to operate once every 1.50 s?
Answer #1

In an charging RC circuit, Voltage is given by:

V = V0*[1 - exp(-t/RC)]

V0 = Initial voltage = 12.0 V

V = Voltage after 1.50 sec = 10.0 V

C = 110 uF = 110*10^-6 F

t = 1.50 sec

So,

V = V0*[1 - exp(-t/RC)]

V/V0 = [1 - exp(-t/RC)]

exp (-t/RC) = 1 - V/V0

-t/RC = ln (1 - V/V0)

R = -t/(C*ln (1 - V/V0))

Using given values:

R = -1.50/(110*10^-6*ln (1 - (10/12)))

R = 7610.6 Ohm

Please Upvote.

Comment below if you've any query.